### v2

# references.bib

Abstract
Road unevenness significantly impacts the safety and comfort of various traffic participants, especially vulnerable road users such as cyclists and wheelchair users. This paper introduces StreetSurfaceVis, a novel dataset comprising 9,122 street-level images collected from a crowdsourcing platform and manually annotated by road surface type and quality. The dataset is intended to train models for comprehensive surface assessments of road networks. Existing open datasets are constrained by limited geospatial coverage and camera setups, typically excluding cycleways and footways. By crafting a heterogeneous dataset, we aim to fill this gap and enable robust models that maintain high accuracy across diverse image sources. However, the frequency distribution of road surface types and qualities is highly imbalanced. We address the challenge of ensuring sufficient images per class while reducing manual annotation by proposing a sampling strategy that incorporates various external label prediction resources. More precisely, we estimate the impact of (1) enriching the image data with OpenStreetMap tags, (2) iterative training and application of a custom surface type classification model, (3) amplifying underrepresented classes through prompt-based classification with GPT-4o or similarity search using image embeddings. We show that utilizing a combination of these strategies effectively reduces manual annotation workload while ensuring sufficient class representation.
Keywords Street-level Imagery, Road Surface Type and Quality, OpenStreetMap, Mapillary
1 Introduction
Road damages have a significant impact on the comfort and safety of all traffic participants and have been identified as a major cause of traffic accidents [kurebwa2019study]. This is particularly true for vulnerable road users such as cyclists [gadsby_understanding_2022, nyberg_road_1996], wheelchair users [pearlman_pedestrian_2013, duvall_development_2013, beale2006mapping] and individuals employing inline skates [lorimer2015beyond], cargo bikes [athanasopoulos2024integrating], scooters [rodier2003unsafe], or strollers. This issue has sparked a large body of research on methods that apply deep learning models to street-level imagery for road surface condition assessment [cao_survey_2020, rahman_comparative_2022, kim_review_2022]. Yet, road damage may not reflect the full range of factors that influence a driver’s experience. For example, the smoothness of sett (regular-shaped cobblestone) is rather determined by the flatness of the utilized stones. Rateke et al. [rateke_road_2019] developed a hierarchical vision-based approach, first predicting the surface type and then employing specific models for each type to classify quality. However, the utilized datasets are too restricted to obtain models that generalize well to new data [rateke_road_2019]. Typical street-level image datasets are commonly collected in good weather conditions, using only a single vehicle and camera setup within a limited geographic boundary, e.g., KITTI [doi:10.1177/0278364913491297], CaRINA [7795529], or Oxford RoboCar [maddern20171]. Cityscapes [cordts_cityscapes_2016] is a dataset of 25,000 images of street scenes recorded in 50 mostly German cities tailored for autonomous driving applications. Perspectives of sidewalks and cycleways are not considered and labels consist of semantic segmentation, while surface type and quality information are unavailable.
We aim to create a street-level image dataset containing a substantial amount of images for each pertinent surface type and quality, fostering the training of robust classification models. We utilize the crowdsourcing platform Mapillary [noauthor_mapillary_2024] to gather images, as these are contributed by individuals from various regions using different devices, camera angles, and modes of transportation, resulting in a heterogeneous dataset. According to data from the crowdsourcing geographic database OpenStreetMap (OSM), distributions of surface type and quality are highly skewed in Germany: For instance, asphalt is the predominant road type, accounting for 47% of tagged road segments, while only about 3% are made of sett. Similarly, for 54% of asphalted roads, the quality is considered ‘good’, while only 1% is rated ‘bad’. Consequently, the difficulty lies in gathering sufficient images for every relevant class without an infeasible manual labeling effort.
This paper introduces StreetSurfaceVis, a new street-level image dataset comprising 9,122 images from the crowdsourcing platform Mapillary, labeled for road surface type and quality by human experts [^1] https://doi.org/10.5281/zenodo.11449977
The referenced repository contains the dataset, a description, a labeling guide, and a datasheet documenting the dataset.
. Additionally, we present and evaluate different strategies for semi-automated annotation to efficiently amplify underrepresented classes in the dataset. These strategies include pre-filtering using OSM tags, iterative training and application of type classification models, prompt-based image classification with GPT-4 models, and similarity-based search using image embeddings.
2 Dataset construction
Image base.
Our dataset is based on images from Mapillary [noauthor_mapillary_2024] limited to the geographical boundaries of Germany. Launched in 2013, the crowdsourcing platform Mapillary provides openly available street-level images. Contributors can use, among others, the Mapillary smartphone app to capture georeferenced image sequences during their trips by, e.g., car, bicycle, or on foot. Thus, the dataset encompasses not only images from roadways but also cycleways and footways. As of January 2024, Mapillary contains about 170 million images in Germany, with more than 50% captured within the last three years. The geographic coverage depends on the contributors within each region, resulting in hundreds of thousands of images for every major city in Germany. As any contributor can upload images, the dataset exhibits varying quality levels, influenced by factors such as the device used, its mounting (e.g., the car dashboard is visible in the frame), or prevailing lighting and weather conditions. Hence, some images are dark, blurred, or contain motifs other than street-level scenes, such as traffic signs, cars, or trees.
Image selection.
To increase the dataset’s heterogeneity, we restrict the number of images for each location and sequence [^2] During recording, the Mapillary app captures images every few seconds, with all images from a trip belonging to the same sequence. per class. This reduces the number of images taken by the same person on one trip and thus increases spatial diversity, camera specifications, environmental conditions, and photographic perspectives. Specifically, we limit the number of images per geographic unit (XYZ style spherical mercator tiles on zoom level 14333We thereby adhere to the same geographic unit as utilized by the Mapillary API for computational feasibility., which is roughly equivalent to $\sim 1.5$ x $1.5km^{2}$ grid cells) to 5 and the number of images per sequence to 10.
For modeling, we recommend using a train-test split where the test data includes geospatially distinct areas, thereby ensuring the model’s ability to generalize to unseen regions is tested. We propose five cities varying in population size and from different regions in Germany for testing, comprising 776 images tagged accordingly.
Labeling scheme.
Our labels for surface type and quality primarily align with the OSM road segment tags surface and smoothness, respectively. While surface [noauthor_osm_2024-1] describes the surface type such as ‘asphalt’, smoothness [noauthor_osm_2024] reflects the physical usability of a road segment for wheeled vehicles, particularly regarding its regularity or flatness [noauthor_osm_2024]. Our labeling scheme includes those classes that are important from a traffic perspective and represent a relevant portion of street types in Germany. This results in the type labels asphalt, concrete, paving stones, sett, and unpaved. For the quality label, we restrict to five of eight proposed levels, ranging from excellent (suitable for rollerblades), good (racing bikes), intermediate (city bikes and wheelchairs), bad (normal cars with reduced velocity) to very bad (cars with high-clearance). The final scheme comprises 18 classes of type and quality combinations, displayed in Table 1, as not all quality labels are suitable for all surface types. See Figure 1 for example images and labels.
(a) paving st. - excel.
strubbl|3044550732526999
(b) paving st. - inter.
carlheinz|789444188806153
(c) unpaved - inter.
macsico|148771344320493
(d) sett - good
hubert87|812903276289693
(e) sett - bad
carlheinz|129178263353193
(f) asphalt - bad
zoegglmeyr|293245499014573
Figure 1: Example images of different surface types and qualities, with Mapillary contributor names and image IDs.
Manual annotation.
After conducting a thorough explorative analysis of Mapillary images, the first three authors developed an annotation guide containing quality level descriptions with example images and underwent self-organized training to manually label surface type and quality. The instructions include labeling the focal road located in the bottom center of the street-level image. In cases where the focus is ambiguous, such as when two parts of the road (e.g., the cycleway and footway) are depicted equally, or when the surface could not be classified due to factors such as snowy roads, blurry images, or non-road images, the image is sorted out. If the surface quality falls between two categories, annotators are directed to select the lower quality level. Annotators are encouraged to consult each other for a second opinion when uncertain. For annotation, we use Labelstudio [LabelStudio], as it allows to preset labels from pre-labeling strategies.
To evaluate inter-rater reliability, 180 heterogeneous images were independently rated by all three annotators. Twelve images were marked for revision by at least one annotator, and a further 49 were discarded by at least one annotator for the above reasons, all of which were excluded for the calculation of inter-rater reliability. Krippendorff’s $\alpha$ [krippendorff_content_2019] for surface type is calculated at 0.96, indicating a high level of agreement. Surface quality, treated as an ordinal scale variable, achieves a Krippendorff’s $\alpha$ of 0.74. While this is generally deemed an acceptable level of agreement, it reflects the fluid class transitions of quality.
An overview of the final annotated dataset is given in Table 1.
Table 1: Final dataset size by type-quality class.
excellent good intermediate bad very bad
asphalt 971 1,696 821 246 -
concrete 314 350 250 58 -
paving stones 385 1,063 519 70 -
sett - 129 694 540 -
unpaved - - 326 387 303
3 Annotation strategies
Using the frequency distribution of OSM tags as a baseline estimate [^4] The OSM distribution is likely an overestimation of the frequency of underrepresented classes, as main roads with good quality are typically more frequented, and thus presumably have more images., manual labeling of randomly sampled images would be highly inefficient. For example, to obtain 7 images of asphalt-bad, 1,000 images would need to be checked (cf. Table  2). Thus, our goal is to employ pre-selection strategies that yield samples for manual labeling where underrepresented classes have a higher frequency than indicated by the OSM baseline. We evaluate three strategies: (1) enriching the image dataset with OSM tags; (2) iterative training and application of a model classifying surface type; and (3) amplifying underrepresented type-quality classes prompting with GPT-4o or through a similarity search using image embeddings. Table  2 presents the improvements through the first two sub-strategies in terms of precision*100, based on 19,747 images processed according to the strategy shown in Figure 2 (marked in green and yellow).
Table 2: Precision*100 of (1) OSM pre-label and (2) type pseudo-label sub-strategies based on 19,747 images vs. OSM frequency distribution as a baseline.
class OSM distr. OSM pre-l. OSM pre-l.
+ type pseudo-l.
asphalt-excellent 20.31 27.83 43.46
asphalt-good 29.29 28.54 50.17
asphalt-intermediate 4.05 14.73 20.31
asphalt-bad 0.69 5.64 8.80
concrete-excellent 0.36 4.23 28.48
concrete-good 0.94 23.32 55.03
concrete-intermediate 0.78 22.34 44.58
concrete-bad 0.23 9.12 20.51
paving stones-excellent 2.98 10.87 30.27
paving stones-good 10.62 15.35 46.36
paving stones-intermediate 1.94 3.05 9.95
paving stones-bad 0.22 1.00 3.83
sett-good 0.56 2.19 12.92
sett-intermediate 1.97 19.37 47.80
sett-bad 1.26 11.92 46.65
unpaved-intermediate 5.75 35.09 60.87
unpaved-bad 7.67 28.61 44.68
unpaved-very bad 7.38 18.70 26.89
Figure 2: Proposed strategy for selecting, pre-labeling, and annotating the training dataset.
Pre-labeling via OSM tags
OSM  [noauthor_openstreetmap_2024] contains surface tags for 49% and smoothness tags for 16% of road segments in Germany, as of January 2024. We incorporate this information by spatially intersecting with geolocations of Mapillary images [^5] For computational feasibility, we refrain from intersecting all Mapillary images and use a sample of Mercantile tiles where the desired classes occur particularly frequently. and assigning the surface and smoothness labels of the closest OSM road segment within a maximum distance of two meters. To eliminate ambiguous street intersections, we cut off 10% of the start and end of each road segment beforehand.
We test the accuracy of this approach based on a sample of 100 images for each of the 18 classes in terms of OSM labels (Batch 1 in Figure  2). OSM labels and manual annotation agree on 69% of surface type labels, and among these, the quality is correct for 55% of the images. However, the errors are not equally distributed: while 94% of asphalt images are correctly pre-labeled, only 54% of paving stones and sett are correct, both being underrepresented surface types. This strategy thus increases the precision for each class substantially, albeit with varying impact. For example, to obtain 7 images from the underrepresented class asphalt-bad, around 100 instead of 1,000 images need to be reviewed (cf. Table  2). We refer to the labels resulting from this strategy as OSM pre-labels.
Incorrect type labels mainly result from mixing up adjacent road parts, as OSM sometimes lacks separate geometries for cycleways and footways. Differences in quality labels are likely due to varying subjective assessments by OSM contributors. Additionally, GPS inaccuracies as well as time differences between image capturing and road segment tagging in OSM are plausible sources of discrepancies, for both type and quality.
Pseudo-labeling with a type classifier
To further increase pre-selection precision, we iteratively train a classification model to predict the type and use its predictions as pseudo-labels. Since type is easier to classify than quality, a smaller amount of data should be sufficient to obtain valuable pseudo-labels. More precisely, we fine-tune EfficientNetV2-S [tan2021efficientnetv2], pre-trained on ImageNet, on the first annotated batch to predict the type. The model is then applied to the next batch of images. All images where the prediction matches the OSM pre-label are selected for manual annotation; this combination of labels achieves an average precision of 95% for the surface type, with the lowest precision of 89% for paving stones over all batches. To reduce bias towards easy-to-classify examples, a random sample of 10% from the excluded images is manually annotated. In the next iteration, the training set is extended with the manually annotated images from the previous round. The batch composition is also adjusted towards underrepresented classes, aiming for 300-400 images per class. This procedure leads to significant improvements for certain classes. For example, paving stones-excellent achieves a precision of 30% with a total of 342 images, while asphalt-good shows an excess of images totaling 1,334. However, the precision remains very low for some classes, e.g., less than 4% (cf. Table  2) for paving stones-bad with only 30 images gathered at this point. Thus, continuing this procedure would be infeasible.
Expanding underrepresented classes through prompting and similarity search
We evaluate two approaches to efficiently enlarge classes that remain underrepresented. Our first approach uses prompt-based image classification with OpenAI’s GPT-4V [noauthor_gptv_system_card_2024] and GPT-4o [noauthor_hello_2024] models, which can generate textual output from image-text input. Previous studies have demonstrated the potential of GPT-4V for automated image labeling in various application domains, including street intersections [hwang_is_2024] and traffic scenery [wen_road_2023]. GPT-4o, released in May 2024, is expected to be similarly effective at half the cost (at the time of release). As an alternative, we explore similarity search using image embeddings [coleman_similarity_2022] from OpenAI’s CLIP  [radford_learning_2021], DINOv2 [oquab_dinov2_2023], and our fine-tuned EfficientNet-based type classifier. Specifically, annotated images from the class of interest are used as the query, and all images with a cosine similarity score above a certain threshold are pre-labeled as class members.
For both approaches, we restrict the search space using the previously described strategy combining OSM tags and type classification. Due to cost and time constraints, we limit the experiments to three underrepresented classes. Using a validation dataset of randomly sampled manually annotated images (50 for the classes asphalt-bad and paving stones - intermediate and 30 for paving stones - bad), we systematically evaluate base models and hyperparameters of both approaches. Note that in both cases, a higher precision implies a lower estimated effort in human post-annotation, while a higher recall correlates with a larger class increase and, in the prompting scenario, with a lower monetary cost.
Prompting configurations include two different image cropping styles, varying levels of detail in class definitions, zero-shot versus one-shot prediction, and processing a batch versus one image per prompt. While cropping shows only minimal effect, one-shot outperforms the zero-shot setting, albeit with a larger monetary cost, resulting in a similar price per hit. GPT-4o achieves notably better results than GPT-4V for all classes. The best configuration in terms of F1 score and cost-effectiveness consists of GPT-4o, lower half-center cropping, shortened definitions, one-shot prediction, and one request per prompt (see Appendix B for the utilized final prompt). To decide on the similarity threshold, we compute the optimal ROC cut-off value for each embedding model and class. All three embeddings give similar and reasonable results for the two larger classes with optimal cut-off values around 0.85 for CLIP and 0.6 for the other two models. The type-classifier slightly outperforms DINOv2 and CLIP on average and is therefore selected for the following experiment. The smallest class paving stones - bad remains difficult for all three models, e.g., our embedding returns nearly all images given the optimal threshold.
The models are then applied in their optimal configurations to a dataset compiled as follows: around 20M Mapillary images are used for image selection as described in Section 2 and pre-labeled using OSM tags and type pseudo-labels. From the remaining data, up to 1,000 images per pre-label are drawn at random. Note that for paving stones-bad only 210 images are available. As shown in Table 3, both approaches yield a substantial improvement. However, prompting with GPT-4o achieves substantially higher precision than the similarity search, with values between 40% and 65%.
Table 3: Evaluation of prompt-based classification (GPT-4o) and similarity search (SimS) on three underrepresented pre-label classes, in terms of precision*100 (recall*100), vs. OSM pre-label + type pseudo-label as a baseline.
class OSM pre-l.
+ type pseudo-l.
OSM pre-l.
+ type pseudo-l.
+ GPT-4o
OSM pre-l.
+ type pseudo-l.
+ SimS
asphalt-bad 10.06 39.46 (81.11) 20.38  (47.78)
paving st.-bad 10.00 64.71 (52.38) 10.29 (100.00)
paving st.-inter. 22.75 40.29 (83.03) 30.74  (47.88)
While the described strategy has shown to have a notable impact (cf. Table 3), it substantially limits the search space (for both GPT-4o and the similarity search) due to the sparsity of OSM tags. Moreover, the reliance on OSM tags yields a potential selection bias as the distribution can be assumed to depend on factors such as urbanity and OSM community. To estimate the efficacy of prompting with GPT-4o without OSM-based filtering, we apply only the type classification model to a random sample of 20,000 Mapillary images to obtain pseudo-labels and thus define the search space. From the resulting 15,100 images classified as asphalt, we prompt GPT-4o with a random sample of 2,000 images, as well as all 712 images pseudo-labeled as paving stones. The results show a reduced precision of 18%-31%, thus higher manual labeling effort. Nevertheless, there remains a relevant increase to the OSM pre-label baseline, providing a viable method if the OSM tag pre-labeled search space is exhausted. However, costs substantially rise: with about \$0.01 per GPT-4o prompt, obtaining a correctly labeled image without pre-labeling via OSM costs \$0.12 to \$3.45, depending on the class, compared to \$0.07 and \$0.18 in the previous experiment.
4 Conclusion and discussion
We presented StreetSurfaceVis, a new street-level imagery dataset labeled according to surface type and quality. To avoid inefficient manual labeling work, we tested different strategies, in particular, to increase the pre-selection of under-represented classes for further manual annotation. As the most feasible strategy, we identified a combination of intersecting OSM tags, predicting surface type with a classification model, and prompting with GPT-4o. We hope that our results will be useful for other researchers and practitioners to efficiently create similar datasets, e.g., for other regions.
Generally, performance increases of GPT-4o and similar models are to be expected in the future, further amplifying the viability of this approach. To reduce dependency on OpenAI and monetary cost, future work should evaluate open-source alternatives. Similarity search is more efficient in computational and monetary terms, therefore currently remaining a viable alternative despite inferior results. The results varied between classes and showed the worst performance for paving stones - bad, where there was no improvement to the baseline. Further experiments, especially with more images for the smallest class, as well as advancements, could be explored, e.g., by incorporating different clustering strategies [vo_automatic_2024].
We envision extensions of the dataset to include road-type information as well as unsuitable images, such as blurred ones or those without a road in focus. Including these will enable training models to automatically check image suitability, thus improving classification robustness and enhancing the GPT-4o prompting strategy.
[…]
Appendix A Labeling scheme
A.1 Asphalt
Quality Description
excellent As good as new asphalt, on which a skateboard or rollerblades will have no problem.
good Asphalt showing the first signs of wear, such as narrow, smaller than 1.5 cm cracks, or wider cracks filled up with tar, shallow dents in which rainwater may collect, which may cause trouble for rollerblades but not for racing bikes.
intermediate Asphalt roads that show signs of maintenance, such as patches of repaired surface, wider cracks larger than 2cm. Asphalt sidewalks may contain potholes, but these are small, shallow (<3cm deep) and can be easily avoided. Asphalt driving lanes show damage due to subsidence (depressions of a scale >50 cm) or heavy traffic (shallow ruts in asphalt caused by trucks in summer). This means that the road can be used by normal city bikes, wheelchairs and sports cars, but not by a racing bike.
bad Damaged asphalt roads that show clear signs of maintenance: This might include potholes, some of them quite deep, which might decrease the average speed of cars. However, it isn’t so rough that ground clearance becomes a problem. Meaning that the street causes trouble to normal city bikes but not a trekking bike and a car.
A.2 Concrete
Quality Description
excellent As good as new concrete, on which a skateboard or rollerblades will have no problem.
good Concrete road showing the first signs of wear, such as narrow, smaller than 1.5 cm cracks, or wider cracks filled up with tar, shallow dents in which rainwater may collect, which may cause trouble for rollerblades but not for racing bikes.
intermediate Concrete roads that show signs of maintenance, such as patches of repaired surface, wider cracks larger than 2cm. Concrete sidewalks may contain potholes, but these are small, shallow (<3cm deep) and can be easily avoided. Concrete driving lanes show damage due to subsidence (depressions of a scale >50 cm) or heavy traffic (shallow ruts in concrete caused by trucks in summer). This means that the road can be used by normal city bikes, wheelchairs and sports cars, but not by a racing bike.
bad Heavily damaged concrete roads that badly need maintenance: many potholes, some of them quite deep. The average speed of cars is less than 50% of what it would be on a smooth road. However, it isn’t so rough that ground clearance becomes a problem. Meaning that the street causes trouble to normal city bikes but not a trekking bike and a car.
A.3 Paving stones
Quality Description
excellent Newly installed and regularly laid paving stones that show no signs of wear. Gaps may be visible, but they are small and uniform and do not significantly affect the driving experience.
good Paving stones showing first signs of wear or newly installed stones with visible but uniform gaps between them. While still suitable for most activities, these surfaces may pose minor challenges for rollerblades and skateboards but remain navigable for racing bikes.
intermediate Characterized by paving stones exhibiting multiple signs of wear, such as shifted heights, potholes, or cracks. This grade allows for the comfortable passage of normal city bikes and standard vehicles but may prove challenging for racing bikes.
bad Heavily uneven or damaged paving stones in dire need of maintenance, featuring significant height disparities and numerous deep potholes. While ground clearance remains sufficient for most vehicles, the surface severely impedes travel, particularly for standard city bikes.
A.4 Sett
Quality Description
good The best sett roads with flattened stones and gaps that are at most small or filled up. The surface might cause trouble for rollerblades but not for racing bikes.
intermediate The surfaces of the sett stones are not completely flat, or there may be slightly larger gaps between the stones, this causes problems for racing bikes and slows down city bikes and cars.
bad Sett stones with large and possibly uneven gaps or uneven stone surfaces or damaged stones, resulting in an overall bumpy surface: This results in a highly uncomfortable driving experience for city bikes. The average speed of cars is less than 50% of what it would be on a smooth road. However, it isn’t so rough that ground clearance becomes a problem.
A.5 Unpaved
Quality Description
intermediate The best unpaved roads that have a compacted surface. This grade allows for the passage of normal city bikes and standard vehicles but may prove challenging for racing bikes.
bad Unpaved roads that do not have a smooth and compacted surface but ones that can still be used with a car or trekking bike but not with a city bike. This category also includes hiking paths, which are too narrow for cars.
very_bad Unpaved roads with potholes, ruts or generally a highly uneven surface not safely passable with a regular passenger car but still passable with an average SUV with higher ground clearance. This category also includes highly uneven hiking paths, which are too narrow for cars.
Appendix B GPT 4-o Prompt
The following shows the utilized prompt for one-shot asphalt image prediction using GPT 4-o. Prompts for other surface types are adjusted according to type-specific labeling schemes (see Appendix  A).
    "role": "system",
    "content": [
            "type": "text",
            "text": "You are a data annotation expert trained to
            classify the quality level of road surfaces in images."
    "role": "user",
    "content": [
            "type": "text",
            "text": f"""
                    You need to determine the quality level of the road
                    surface depicted in the image, following this defined scale:
                    Asphalt surfaces are graded from excellent to bad
                    according to the following scale:
                    1) excellent: As good as new asphalt, on which a
                    skateboard or rollerblades will have no problem.
                    2) good: Asphalt showing the first signs of wear, such as
                    narrow, smaller than 1.5 cm cracks, or wider cracks filled
                    up with tar, shallow dents in which rainwater may collect,
                    which may cause trouble for rollerblades but not for
                    racing bikes.
                    3) intermediate: Asphalt roads that shows signs of
                    maintenance, such as patches of repaired surface, wider
                    cracks larger than 2cm. Asphalt sidewalks may contain
                    potholes, but these are small, shallow (<3cm deep) and can
                    be easily avoided, asphalt driving lanes shows damage due
                    to subsidence (depressions of a scale >50 cm) or heavy
                    traffic (shallow ruts in asphalt caused by trucks in
                    summer). This means that the road can be used by normal
                    city bikes, wheelchairs and sports cars, but not by a
                    racing bike.
                    4) bad: Damaged asphalt roads that show clear signs of
                    maintenance: This might include potholes, some of them
                    quite deep, which might decrease the average speed of
                    cars.  However, it isnt so rough that ground clearance
                    becomes a problem. Meaning that the street causes trouble
                    to normal city bike but not a trekking bike and a car.
                    Please adhere to the following instructions:
                    1) Step 1: If you detect multiple surface types, only
                    consider the path, driving lane, cycleway or sidewalk in
                    the focus area.
                    2) Step 2: Check if the road surface is worn off and if
                    you can find any damages, like cracks.
                    3) Step 3: Check the quantity and the size of the damages.
                    3) Step 4: Then decide if you could ride on the surface
                    with  a skateboard, rollerblades, racing bikes, city bike,
                    or a normal car.
                    4) Step 5: If you detect characteristics of two classes,
                    choose the worse class.
                    How would you rate this image using one of the four
                    options of the defined scale:
                    1) excellent
                    2) good
                    3) intermediate
                    4) bad
                    Provide your rating in one word disregarding the bullet
                    point numbers and brackets as a string using the four
                    levels of the scale provided.
                    Make sure you have the same number of image urls as input
                    as you have output values.
                    Do not provide any additional explanations for your
                    rating; focus solely on the road surface quality."""
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{excellent_encoded_image}"
            "type": "text",
            "text": "This was an example for ’excellent’"
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{good_encoded_image}"
            "type": "text",
            "text": "This was an example for ’good’"
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{intermediate_encoded_image}"
            "type": "text",
            "text": "This was an example for ’intermediate’"
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{bad_encoded_image}"
            "type": "text",
            "text": "This was an example for ’bad’"
            "type": "text",
            "text": "Please decide now in one word which category the
            following picture belongs to as instructed in the beginning of the
            prompt. Then compare this image to the previous ones and decide
            whether this category is correct."
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{image}"